why refractive index and wavelenght is inversely proportional

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
As we all know that
Refractive index = speed of light in vacuum / speed of light in medium
  • = c/v
v =ƛ/t
  • = c / ƛ/t
  • = ct/ ƛ
µ = 1/ƛ
From  the above expression
µ is inversely proportional to wavelength. 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)